OpenStudy (anonymous):

The sum of the reciprocals of two consecutive even integers is 9/40 Write an equation that can be used to find the two integers. What are the integers?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Let x = first integer x+2 = next consecutive even integer 1/x=reciprocal of first integer 1(/x+2)= reciprocal of next consecutive even integer Write the equation that says the sum of the reciprocals = 9/40

OpenStudy (anonymous):

So 1(/x+2)=9/40?

OpenStudy (mertsj):

What is the reciprocal of the first integer?

OpenStudy (anonymous):

1/x

OpenStudy (mertsj):

What is the reciprocal of the next consecutive even integer?

OpenStudy (anonymous):

x+2

OpenStudy (mertsj):

Reread my first post. What is the RECIPROCAL of the next consecutive even integer?

OpenStudy (anonymous):

1(/x+2)

OpenStudy (mertsj):

How do you find a sum?

OpenStudy (anonymous):

By multiplying?

OpenStudy (mertsj):

A sum is the answer to an addition problem.

OpenStudy (mertsj):

What is the sum of the two RECIPROCALS?

OpenStudy (anonymous):

How could you add?

OpenStudy (mertsj):

The sum of 2 and 3 is 2+3 The sum of x and y is x+y The sum of 1/x and 1/(x+2) is ?????

OpenStudy (anonymous):

OH! 1/x +1/(x+2) 2/2x+3?

OpenStudy (mertsj):

\[\frac{1}{x}+\frac{1}{x+2}=\frac{9}{40}\]

OpenStudy (anonymous):

Sorry, im just so confused.

OpenStudy (mertsj):

Solve that equation. Multiply both sides by the common denominator which is 40x(x+2)

OpenStudy (anonymous):

80 x (x+2)+40 (x+2)+1/x = 9 x (x+2)?

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

x=-10/9,8?

OpenStudy (anonymous):

..

OpenStudy (anonymous):

Mertsj, are you there?

OpenStudy (mertsj):

Sorry. I was called away.

OpenStudy (mertsj):